OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 45 when x = 9, find the value of f(x) when x = 3. A. 5 B. 9 C. 15 D. 27

OpenStudy (anonymous):

hint : varies directly means when x increases then f(x) increases and when x decreases then f(x) decreases

OpenStudy (anonymous):

so 27?

OpenStudy (anonymous):

i.e. 45/9 = f(x)/3 f(x) = 3*45/9 = ?

OpenStudy (anonymous):

it's not 27

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

excellent!

OpenStudy (anonymous):

@chaser71

OpenStudy (anonymous):

This is a direct proportions problem. 45 / 9 = ? / 3 What should be the number in the place of the question mark make sense? :D

OpenStudy (anonymous):

oh!